# Learning to Generate Instruction Tuning Datasets for Zero-Shot Task Adaptation

### v1

@@ 5.1 Experiment Setup
In this work, we consider three target tasks: yes-no question answering (YNQA), extractive question answering (ExQA), and natural language inference (NLI). Table 1 shows the seven datasets along with the number of unannotated texts across three task types in our experiments. For yes-no question answering, we choose PubMedQA (Jin et al., 2019) and Privacy Policy QA (Ravichander et al., 2019). For extractive question answering, we choose the SquadShifts dataset (Miller et al., 2020) which includes splits for the New York Times (NYT), Amazon, and Reddit. Finally, for the NLI task, we choose Contract-NLI (Koreeda and Manning, 2021) and Vitamin C (Schuster et al., 2021). We provide additional details in Appendix A.
@@
In our experiments, we focus on tasks that require a two-step task generation process, i.e., first, we need to generate a question or a hypothesis before generating the answer. Prior work generates synthetic tasks like summarization that do not warrant a specialized task generation model (Yehudai et al., 2024). They also generate instructions (Li et al., 2023; Köksal et al., 2023) for long-form conversational datasets where the solution to the instruction is the unannotated text. While these long-form synthetic tasks are useful for applications such as code generation, domains like biomedical and legal that we consider might benefit more from traditional predictive rather than generative tasks (Miller, 2024).
@@
We consider two key baselines: zero-shot and self supervised baseline. For the zero-shot baseline, we simply prompt the model and run the evaluation without using any of the unannotated text from the target task (None). For the self supervised baseline, we use task-adaptive pretraining (TAPT) (Gururangan et al., 2020). The learning objective is to continue to the pretraining objective on the unannotated text in the downstream dataset. In our experiments, we use the next word prediction learning objective to fine-tune Mistral-7B and Llama 2 7B models.
@@
Here we describe the process of generating synthetic tasks with Bonito. As described in Section 4, given a task type, we prompt Bonito with the unannotated texts and task types to generate the instruction tuning data. We use nucleus sampling (Holtzman et al., 2020) with a top P value of and a temperature of , and a maximum sequence length of in the vLLM framework (Kwon et al., 2023).
The generated tasks are post-processed into a standardized instruction-response format for instruction tuning. In each generation, we replace {context} with the actual unannotated text If the generated output is not parsable due to missing <|pipe|>, we filter them out.
@@
We consider adapting two pretrained large language models: Mistral-7B (Jiang et al., 2023) and Llama 2 7B (Touvron et al., 2023). They are decoder language models trained with the next word prediction objective on trillions of tokens. Both these models are 7 billion parameters in size with slightly different architectures optimized for sequence length and inference. For more details, see Touvron et al. (2023) and Jiang et al. (2023).
We also consider a more practical setting where we further adapt instruction tuned model to the target task. We first consider an off-the-shelf instruction tuned model: Mistral-7B-Instruct-v0.2. This model based on Mistral-7B achieves comparable performance to Llama 2 13B Chat on the MT-Bench (Zheng et al., 2023). In addition, we train Mistral-7B and Llama 2 models on the T0 split from the P3 dataset (Bach et al., 2022) and adapt them to the target tasks. We refer to these models as Mistral-7BP3 and Llama 2P3. For the instruction tuning details, see Appendix E.2
@@
We fine-tune the language models on the supervision sources, TAPT, and Bonito, using Q-LoRA (Dettmers et al., 2023). When further adapting Mistral-7BP3 and Llama 2 7BP3, we fine-tune the same Q-LoRA adapter on the supervision sources instead of merging and reinitializing the adapters. We train all the models for 1 epoch. If the dataset size is greater than 160,000 examples, then we train for 10,000 steps. To avoid additional hyperparameter tuning, we use the same hyperparameter values from Dettmers et al. (2023). Depending on the number of steps and the dataset, training on four GPUs takes between 25 minutes to 17 hours. For more additional details, see Appendix E.5.
@@
We evaluate the performance of the models on the test splits of the target datasets. To prevent “prompt hacking”, following Sanh et al. (2022), we first write five prompt templates for target datasets and then benchmark the model performance. See Appendix H for all the prompts used in our experiments. We follow standard evaluation practices and report the F1 score for all the datasets. Following Radford et al. (2019), to evaluate models on yes-no question answering and NLI, we use ranked classification, i.e., generate the loglikelihood of all the choices and choose the sequence with the highest loglikelihood as the prediction. Following Rajpurkar et al. (2016), we evaluate models on extractive question answering by computing the SQuAD F1 score on the generated output. During evaluation, we use greedy decoding to generate the output from the model and then calculate the SQuAD F1 score for the dataset.
@@ 5.3 Adapting Instruction Tuned Models
Table 3 shows that Bonito improves instruction tuned models by an average of 22.1 F1 points whereas TAPT reduces the average performance by 0.8 F1 points. This is because self supervision with TAPT interferes with prior instruction tuning and leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017). In contrast, we find that adapting instruction tuned models with Bonito-generated tasks further improves performance on tasks in specialized domains. We observe that Bonito addresses the task-specific deficiencies and improves the instruction tuned models. For example, we find that Bonito significantly improves Mistral-7B-Instruct-v0.2 performance on extractive question answering as it typically generates chat-like responses for questions. Finally, we find that adapting instruction tuned variants of Mistral-7B and Llama 2 7B achieves a higher F1 score than adapting the pretrained models (see Table 2).
@@ 6.1 Impact of Domain Knowledge
Here we ask a key question: are we improving the language model by learning about the domain or are we distilling instructing tuning data from a stronger to a weaker model? To answer this question, we train task-specialized instruction tuned models and then further train them on synthetic tasks generated from Bonito for the target unannotated texts. We create the task-specialized training dataset by selecting prompts in datasets of the target task type. We train two task-specialized models: Mistral-7B-Instruct-v0.2special and Mistral-7Bspecial. We create meta-templates from the same dataset to train a task-specialized Bonito special. See Appendix E.3 for training details.
Table 4 shows that further training on synthetic instructions can further improve performance which suggests that the model benefits from the unnannotated text from the specialized domain. We find that training on Bonito tasks either slightly improves or matches the performance of task-specialized models on average. When we train on Bonito special tasks, we further improve task-specialized Mistral-7B-Instruct-v0.2 by 0.5 F1 points and Mistral-7B and 2.5 F1 points. We see that the model performance often reduces on extractive QA. We suspect that the model performance has saturated due to the presence of SQuAD in the task-specialized training dataset. To further improve on extractive question answering, we could benefit from having access to a few examples from the target dataset. Finally, we almost always improve performance on Vitamin C and PubMedQA datasets highlighting the importance of training on more task samples (see Section 6.2).
@@ 6.2 Effect of the Training Dataset Size
Here we study the effect of the size of the training dataset. In particular, we study how Mistral-7B performance varies on when trained on different quantities of synthetic instruction tuning data for PubMedQA and Vitamin C. Figure 3 shows that training on more steps typically improves performance. We find that Bonito on PubMedQA reaches the peak performance of 47.1 F1 points after 10,000 steps but the F1 can fluctuate when trained for fewer steps. In contrast, we find that Bonito gets the highest performance of 73.3 F1 points after 2500 points and gradually diminishes the performance to 71.7 F1 points. Finally, we suggest using a validation set, if available, to select the best-performing model checkpoint.
@@ 7 Additional Experiments
In Appendix B, we generate synthetic tasks by prompting Mistral-7B-Instruct-v0.2 and Zephyr-7B-. Our results show that the synthetic tasks improve the average performance of Mistral-7B but decrease significantly when adapting MistralP3. This shows that naively generating synthetic tasks is not sufficient, and we require high-quality synthetic tasks to increase the performance of strong instruction-tuned models.
In Appendix C, we generate synthetic tasks with GPT-4 for Privacy Policy QA, SQuADShifts Reddit, and ContractNLI. Our results show that GPT-4 improves MistralP3 on Privacy Policy QA and ContractNLI but slightly reduces performance on SQuADShifts Reddit.
We analyze the generated tasks and identify common issues in both open-source models and GPT-4, such as the distribution of the label space and “chatty” responses, as potential causes for the drop in performance.
@@ 8 Conclusion
We present Bonito, an open-source model for conditional task generation to convert unannotated texts into instruction tuning datasets. We show that training with synthetic instruction tuning datasets in specialized domains is a strong alternative to self supervision. Our experiments demonstrate that Bonito-generated instructions improve both pretrained and instruction tuned models on zero-shot task adaptation. Overall, Bonito enables practitioners to adapt large language models to tasks on their data without annotations.
@@ Appendix A Datasets
We briefly describe the datasets used in our experiments. We obtain all the datasets from the datasets library (Lhoest et al., 2021). Table 5 shows the statistics for the test sets in the evaluation datasets. For all the datasets, we consider five prompt templates (see Appendix H). Below we include details about the evaluation datasets:
PubMedQA (Jin et al., 2019): The dataset is about biomedical research questions, utilizing context from PubMed abstracts that can be answered with yes, no, or maybe. The original PubMedQA dataset has two settings: reasoning-required and reasoning-free. In this paper, we provide context, the PubMed Abstract, to the model, ensuring that all results reported fall within the reasoning-required setting.
Privacy Policy QA (Ravichander et al., 2019): The dataset consists of paragraphs from privacy policies paired with corresponding questions. The task involves determining the relevance of each question, formatted as a yes-or-no question-answering task. We use the processed test split of Privacy Policy QA from Guha et al. (2023) as the unannotated text.
SquadShifts (Miller et al., 2020): The dataset consists of four new test sets for the SQuAD (Rajpurkar et al., 2016). In this paper, we specifically choose three of them — New York Times articles, Reddit posts, and Amazon product reviews. The dataset serves to assess the model’s reading comprehension ability and is structured as an extractive question-answering task.
ContractNLI (Koreeda and Manning, 2021): The ContractNLI requires that given an excerpt of a contract and an assertion about the legal effect of that excerpt, the model need to determine whether the assertion is supported or unsupported by the excerpt. The dataset is prompted into a natural language inference task.
Vitamin C (Schuster et al., 2021): This dataset focuses on fact verification through factual revisions to Wikipedia pages. Each example consists of an evidence text from Wikipedia and a corresponding fact. The model is asked to indicate whether the fact is supported, refuted, or neutral with respect to the evidence.
@@ B.1 Generating Synthetic Datasets
Here we describe the process of creating synthetic datasets with Mistral-Instruct-v0.2 and Zephyr-. We prompt these models to generate questions or hypotheses for the target unannotated text. Table 6 shows the prompts that we used to generate the tasks. Creating these prompts required a tremendous amount of prompt engineering. We first generate the question or the hypothesis and then generate the answer as these models often ignore multiple instructions in the prompt. We then parse the generated question and the hypothesis and re-prompt the model to generate a response. For question answering tasks, we prepend the question as the prompt followed by the unannotated text to generate the output. For the NLI datasets, we use five prompt templates from the ANLI dataset in Bach et al. (2022) and plug in the hypothesis and the unannotated text as the input to the model to generate the answer. We use the same input and output to adapt the pretrained and instruction tuned models. For all the generations, we use a top-P of 0.95, temperature of 0.5, and maximum token length of 256.
@@ B.2 Results
Table 7 shows results for zero-shot task adaptation with openly available models. We see that both Mistral-7B-Instruct-v0.2 and Zephyr-7B- improve performance over the pretrained Mistral-7B but we find that they severely hurt average performance compared to Mistral-7BP3.
We suspect that the drop in performance is due to issues related to the generated tasks. For extractive question answering, we find that Mistral-7b-Instruct-v0.2 and Zephyr- often generate questions with multiple sub-questions that cannot be easily answered by extracting words from the context. Furthermore, the responses are "chatty," which might not be appropriate for extractive question answering. We also observe that many of the generated questions are often positive, i.e., they usually have "yes" or "true" as the answer. For example, PubMedQA generated by Zephyr- has about 68% of the questions starting with "yes" or "true" as the answer, and about 5% have an answer that starts with "no" or "false." We observe a similar trend with the hypotheses generated for natural language inference datasets. For instance, the Contract NLI dataset generated by Zephyr- shows that about 64% have "yes," "true," or "correct" as the answer, whereas only 1% have "no," "false," or "incorrect" as the answer for the hypothesis.

### v2

@@ 5.1 Experiment Setup
We consider three target tasks: yes-no question answering (YNQA), extractive question answering (ExQA), and natural language inference (NLI). Table 1 shows the seven datasets across three task types and the number of unannotated text in each dataset. We use the unannotated text from the datasets to train the specialized language models. For yes-no question answering, we choose PubMedQA (Jin et al., 2019) and Privacy Policy QA (Ravichander et al., 2019). For extractive question answering, we choose the SquadShifts dataset (Miller et al., 2020) that includes splits for the New York Times (NYT), Amazon, and Reddit. Finally, for the NLI task, we choose Contract-NLI (Koreeda and Manning, 2021) and Vitamin C (Schuster et al., 2021). We provide additional details in Appendix A.
@@
In our experiments, we focus on tasks such as question answering and natural language inference that require us to generate a question or hypothesis and an answer. Prior work generates synthetic data for tasks like summarization that do not warrant a specialized task generation model (Yehudai et al., 2024). Other work focuses on generating instructions (Li et al., 2023; Köksal et al., 2023) for long-form text generation tasks where the solution to the instruction is the unannotated text. While these long-form generative tasks are useful for applications such as code generation, domains like biomedical and legal that we consider might benefit more from traditional predictive tasks (Miller, 2024).
@@
We consider two key baselines: zero-shot and self supervised baseline. For the zero-shot baseline, we prompt the model and run the evaluation without using any of the unannotated text from the target task (None). For the self supervised baseline, we use task-adaptive pretraining (TAPT) (Gururangan et al., 2020). The learning objective is to continue to the pretraining objective on the unannotated text in the downstream dataset. In our experiments, we use the next word prediction learning objective to fine-tune Mistral-7B and Llama 2 7B models.
@@
As described in Section 4.1, we prompt Bonito with the unannotated texts and the target task type to generate the instruction tuning data. We use nucleus sampling (Holtzman et al., 2020) with a top P value of and a temperature of , and a maximum sequence length of in the vLLM framework (Kwon et al., 2023).
@@
We adapt two pretrained large language models: Mistral-7B (Jiang et al., 2023) and Llama 2 7B (Touvron et al., 2023). They are decoder language models trained with the next word prediction objective on trillions of tokens. Both these models have around 7 billion parameters, with slightly different architectures optimized for sequence length and inference. For more details, see Touvron et al. (2023) and Jiang et al. (2023).
We also consider a more practical setting where we further adapt instruction tuned models to the target task. We first consider an off-the-shelf instruction tuned model: Mistral-7B-Instruct-v0.2. This model based on Mistral-7B achieves comparable performance to Llama 2 13B Chat on the MT-Bench (Zheng et al., 2023). In addition, we train Mistral-7B and Llama 2 models on the T0 split from the P3 dataset (Bach et al., 2022) and adapt them to the target tasks. We call these models Mistral-7BP3 and Llama 2P3. For the instruction tuning details, see Appendix F.2
@@
We fine-tune the language models on the supervision sources—TAPT, and Bonito—using Q-LoRA (Dettmers et al., 2023). When further adapting Mistral-7BP3 and Llama 2 7BP3, we fine-tune the same Q-LoRA adapter on the supervision sources instead of merging and reinitializing the adapters. We train all the models for 1 epoch. If the dataset size is greater than 160,000 examples, then we train for 10,000 steps. We use the same hyperparameter values from Dettmers et al. (2023) to avoid additional hyperparameter tuning. Depending on the dataset, training on four GPUs takes 25 minutes to 17 hours. For more additional details, see Appendix F.5.
@@
We evaluate the performance of the models on the test splits of the target datasets (see Table 6 in Appendix A). To prevent “prompt hacking”, following Sanh et al. (2022), we first write five prompt templates for the target datasets and then benchmark the model performance. See Appendix I for all the prompts used in our experiments. We follow standard evaluation practices and report the F1 score for all the datasets. Following Radford et al. (2019), we evaluate yes-no question answering and NLI using ranked classification, i.e., we generate the loglikelihood of all the choices and choose the sequence with the highest loglikelihood as the prediction. Following Rajpurkar et al. (2016), we evaluate models on extractive question answering by computing the SQuAD F1 score on the generated output. During evaluation, we use greedy decoding to generate the output from the model and then calculate the SQuAD F1 score for the dataset.
@@ 5.3 Adapting Instruction Tuned Models
Table 3 shows that Bonito improves instruction tuned models by an average of 22.1 F1 points whereas TAPT reduces the average performance by 0.8 F1 points. This is because self supervision with TAPT interferes with prior instruction tuning and leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017). In contrast, adapting instruction tuned models with Bonito-generated tasks further improves performance in specialized domains. We also observe that Bonito addresses the task-specific deficiencies and improves the instruction tuned models. For example, we find that Bonito significantly improves Mistral-7B-Instruct-v0.2 performance on extractive question answering as it typically generates chat-like responses for questions. Finally, adapting instruction tuned variants of Mistral-7B and Llama 2 7B achieves a higher F1 score than adapting the pretrained models (Table 2).
@@ 6.1 Impact of Domain Knowledge
Here we ask a key question: are we improving the language model by learning about the domain or are we distilling instructing tuning data from a stronger to a weaker model? To answer this question, we train task-specialized instruction tuned models and then further train them on synthetic tasks generated from Bonito for the target unannotated texts. We create the task-specialized training dataset by selecting the instructions in CTGA with the target task type. We train two task-specialized models in the standard instruction-response format: Mistral-7B-Instruct-v0.2special and Mistral-7Bspecial. We also train a task-specialized Bonito special on the same task-specific dataset. See Appendix F.3 for training details.
Table 4 shows that further training on synthetic instructions can improve performance suggesting that the model benefits from the unnannotated text from the specialized domain. We find that training on Bonito tasks either slightly improves or matches the performance of task-specialized models on average. When we train on Bonito special tasks, we further improve task-specialized Mistral-7B-Instruct-v0.2 by 0.5 F1 points and Mistral-7B and 2.5 F1 points. We see that the model performance often reduces on extractive QA. We suspect that the model performance has saturated due to the presence of SQuAD in the task-specialized training dataset. To further improve on extractive question answering, we could benefit from having access to a few examples from the target dataset. Finally, we almost always improve performance on Vitamin C and PubMedQA datasets highlighting the importance of training on more task samples (see Section 6.2).
@@ 6.2 Effect of the Training Dataset Size
Here we study the effect of the size of the training dataset. In particular, we study how Mistral-7B performance varies when trained on different quantities of synthetic instruction tuning data for PubMedQA and Vitamin C. Figure 3 shows that training on more steps typically improves performance. We find that Bonito on PubMedQA reaches the peak performance of 47.1 F1 points after 10,000 steps but the F1 can fluctuate when trained for fewer steps. In contrast, we find that Bonito gets the highest performance of 73.3 F1 points after 2,500 points and gradually diminishes the performance to 71.7 F1 points. Finally, if available, we suggest using a validation set to select the best-performing model checkpoint.
6.3 Human Evaluation: Agreement with GPT-4o
We manually evaluate Bonito tasks by comparing the answers generated by Bonito and GPT-4o.
| Dataset | Match=Yes | Agreement (>=2) | ||
| PubMedQA | 72% | 97% | ||
| 76% | 88% | |||
| ContractNLI | 71% | 99% |
Setup
We sample 100 unique instructions each from Bonito for PubMedQA, SQuADShifts Reddit, and ContractNLI. Next, we prompt GPT-4o with instructions generated by Bonito to produce an answer. We prefix the instructions with a simple format prompt to produce answers in the desired format with GPT-4o. Finally, we ask humans if GPT-4o and Bonito produce the same (including paraphrased) answers for the instructions. For reproducibility, we use GPT-4o-2024-05-13. We separately ask the first three authors of the paper to compare the responses from both models. We choose the final agreement if two or more annotators agree on either a match or no match.
Results
Table 5 shows that Bonito and GPT-4o produce the same answer for Bonito tasks 71 to 76 percent of the time across three datasets, with high inter-annotator agreement. Each dataset reveals different patterns of disagreement. In PubMedQA, GPT-4o generates the response as “unanswerable” when the answer is not present in the passage, whereas Bonito produces either “yes” or “no”, or “true” or “false”. In SQuADShifts Reddit, Bonito almost always extracts the answer from the paragraph, whereas GPT-4o can generate answers with additional text that may not be present in the paragraph. In ContractNLI, both models can produce plausible answers. In one example, Bonito generates the question, ‘Does this imply that “This document is confidential information”? Yes, no, or maybe?’. Bonito answers “yes”, whereas GPT-4o produces the answer “maybe”. In such cases, we annotate the responses as no match, reducing the agreement.
Our analysis shows that Bonito generates high-quality tasks with accurate answers. However, there is still room to improve the quality of the tasks. Improving the task quality could further increase the downstream model performance. Therefore, we believe that research on conditional task generation is an important direction for future work.
@@ 7 Additional Experiments
In Appendix B, we generate synthetic tasks by prompting Mistral-7B-Instruct-v0.2 and Zephyr-7B-. Our results show that the synthetic tasks from Mistral-7B-Instruct-v0.2 and Zephyr-7B- improve the average performance of Mistral-7B but decrease significantly when adapting MistralP3. This indicates that we require high-quality synthetic tasks to increase the performance of strong instruction-tuned models. In Appendix C, we generate synthetic tasks with GPT-4 for Privacy Policy QA, SQuADShifts Reddit, and ContractNLI. Our results show that GPT-4 improves MistralP3 on Privacy Policy QA and ContractNLI but slightly reduces performance on SQuADShifts Reddit. Finally, we analyze the generated tasks and identify common issues with both open-source models and GPT-4, such as the distribution of the label space and “chatty” responses, which potentially lead to the drop in performance.
@@ 8 Conclusion
We present Bonito, an open-source model for conditional task generation that converts unannotated text into instruction tuning datasets. We show that training with synthetic instruction tuning datasets in specialized domains is a strong alternative to self supervision. Our experiments demonstrate that Bonito-generated instructions improve pretrained and instruction tuned models on zero-shot task adaptation. Overall, Bonito enables practitioners to adapt large language models to tasks on their data without annotations.
@@ Appendix A Datasets
We briefly describe the datasets used in our experiments. We get all the datasets from the datasets library (Lhoest et al., 2021). For all the datasets, we consider five prompt templates (see Appendix I). Table 6 shows the statistics for the test splits in the evaluation datasets. Below we include details about the evaluation datasets:
PubMedQA (Jin et al., 2019): The dataset asks questions about PubMed abstracts that can be answered with yes, no, or maybe. We use the abstracts without the questions as unannotated text for adaptation. During the evaluation, we provide the PubMed abstract along with the question to the model.
Privacy Policy QA (Ravichander et al., 2019): The dataset consists of paragraphs from privacy policies with corresponding questions. The task involves determining the relevance of each question, formatted as a yes-no question-answering task. We use the processed test split in Privacy Policy QA from Guha et al. (2023) as unannotated text.
SquadShifts (Miller et al., 2020): The dataset is designed to test the robustness of extractive question answering models. We use three of the four test sets in our work — New York Times articles, Reddit posts, and Amazon product reviews. During training, we use the articles or context from the test set without the questions and generate extractive question answering tasks with Bonito. During evaluation, we evaluate the same test set with the questions in the dataset.
ContractNLI (Koreeda and Manning, 2021): ContractNLI is a natural language inference task to aid contract review. Given a hypothesis about a clause in a contract, the model predicts if the hypothesis is supported, refuted, or not mentioned.
Vitamin C (Schuster et al., 2021): This dataset focuses on fact verification in Wikipedia framed as a natural language inference task. Each example consists of an evidence text from Wikipedia and a corresponding fact. The model is asked to indicate whether the fact is supported, refuted, or neutral.
@@ B.1 Generating Synthetic Datasets
Here we describe the process of creating synthetic datasets with Mistral-Instruct-v0.2 and Zephyr-. We prompt these models to generate questions or hypotheses for the target unannotated text. Table 7 shows the prompts we used to generate the tasks. Creating these prompts required a tremendous amount of prompt engineering as they struggled to follow the prompt format (Xia et al., 2024). We first generate the question or the hypothesis and then re-prompt the model to produce the answer. For question answering tasks, we prepend the question as the prompt followed by the unannotated text to generate the output. For the NLI datasets, we use five prompt templates from the ANLI dataset in Bach et al. (2022) and plug in the hypothesis and the unannotated text as the input to the model to generate the answer. We use the same input and output to adapt the pretrained and instruction tuned models. For all the generations, we use a top-P of 0.95, temperature of 0.5, and maximum token length of 256.
@@ B.2 Results
Table 8 shows results for zero-shot task adaptation with openly available models. We see that both Mistral-7B-Instruct-v0.2 and Zephyr-7B- improve performance over the pretrained Mistral-7B but find that they severely hurt average performance compared to Mistral-7BP3.
We suspect that the drop in performance is due to issues related to the generated tasks. For extractive question answering, we find that Mistral-7b-Instruct-v0.2 and Zephyr- often generate questions with multiple sub-questions that cannot be easily answered by extracting words from the context. Furthermore, the responses are “chatty” which might not be appropriate for extractive question answering. We also observe that the generated questions are often “positive”, i.e., they usually have “yes” or “true” as the answer. For example, 68% of the questions generated by Zephyr- for PubMedQA have answers starting with “yes” or “true” but only 5% of the questions have answers that start with “no” or “false”. We observe a similar “positive” bias in the hypotheses generated for natural language inference datasets.